OpenStudy (anonymous):

(Newton's Law of Motion) A mass of 8 kg and another of 12 kg are suspended by a string on either side of a frictionless pulley. Find the acceleration of each mass.

OpenStudy (anonymous):

Hi first you ought right 2nd law of newtonian for each masses and discard "T" to find a\[T-m _{1}g=m_{1}a\]\[m _{2}g-T=m _{2}a\] finally you find:\[a=(m _{1}-m _{2})/(m _{1}+m _{2})g\] . i yield it: a=(g/5)